OpenStudy (anonymous):

Find the derivative of y= x arc cos x - (1-x^2)^(1/2)

OpenStudy (anonymous):

use product rule and chain rule

OpenStudy (amistre64):

can you clean up the equation any?

OpenStudy (anonymous):

think its arccos(x), is the final answer

OpenStudy (anonymous):

i have seen questions like this before

OpenStudy (anonymous):

Does the following look OK? \[-\frac{2 x}{\sqrt{1-x^2}}+\text{ArcCos}[x] \]

OpenStudy (anonymous):

"differentiate this ..... hence find integral of arccos(x)

OpenStudy (anonymous):

yeh that looks correct ^

OpenStudy (anonymous):

then a very common question to have after this would be to find integral of arccos(x)

OpenStudy (anonymous):

\[\int\limits \left(-\frac{2 x}{\sqrt{1-x^2}}+\text{ArcCos}[x]\right) \, dx=\sqrt{1-x^2}+x \text{ArcCos}[x] \]